OpenStudy (anonymous):

What is the exact value of the expression 7√8+2√12-3√2 Show your work.

OpenStudy (anonymous):

7√8 + 2√12 - 3√2 = 7√4√2 + 2√4√3 - 3√2 =14√2 + 8√3 - 3√2 =11√2 + 8√3